OpenStudy (toxicsugar22):

Andy drives 300 miles to Lake Tahoe at 70 miles per hour and returns home at 50 mph. What is his average speed for the round trip. (It is not 6o miles per hour.) Write expressions for the time it takes for each leg of the trip if Andy drives a distance d, at speed r base 1 and returns at speed r base 236 minutes ago

OpenStudy (dumbcow):

average speed = total distance/total time = 600/[300/70 + 300/50] = 600/(30/7 + 6) =600/(72/7) =4200/72 =58.33 mph

OpenStudy (toxicsugar22):

thanks

OpenStudy (toxicsugar22):

can u help me find the time trip to Tahoe

OpenStudy (toxicsugar22):

can u help me

OpenStudy (toxicsugar22):

will u help me

OpenStudy (dumbcow):

sorry my computer is being weird the time = distance/speed t1 = 300/70 = 4.286 t2 = 300/50 = 6

OpenStudy (toxicsugar22):

ok and how do u write 58 as a simple fraction